Ifthe error in the measurement of mass of a body is 3% and in tbe measuremenrt of velocity is 2% what is the maximum possible error in calculation of kinetic energy

Grade:11

1 Answers

Arun
6 years ago
I will be using ‘d’ for delta
K.E. = (1/2) m*v^2
d(KE)/KE = dm/m + 2 dv/v
100*d(KE)/KE = 100*dm/m + 2*100* dv/v
% chenge in KE = 3+ 2*2
 = 7%